OpenStudy (anonymous):

Suppose that a sphere with radius 5a has the same volume as a cone of radius 3a. What is the height of the cone? Give your answer in terms of a. Use the formulas and for the volumes of a sphere and cone respectively.

OpenStudy (anonymous):

\[\frac{ 4 \Pi 5a *5a*5a}{ 3 } = \frac{ 1\Pi 3a*3a * h }{ 3 }\] Find h..